# Real-Time Stochastic Terrain Mapping and Processing for Autonomous Safe Landing

### v1

@@ 1 Introduction
This paper presents a novel real-time stochastic hazard detection algorithm capable of handling topographic uncertainty arising from sparse terrain measurement. First, we develop a novel real-time Gaussian DEM construction algorithm that utilizes Delaunay triangulation and a localized Gaussian Random Field (GRF) model for efficient evaluation. Next, we introduce an innovative real-time stochastic hazard detection algorithm that converts Gaussian DEMs into stochastic safety maps. A fast and provably conservative HD algorithm is developed for standard DEM input, which is then extended to accommodate Gaussian DEM input. Finally, the effectiveness of these proposed methodologies is demonstrated through a combination of real and synthetic terrain data.
@@ 1.1 Background
Historically, the deployment of planetary landers has been constrained by the absence of onboard hazard detection and avoidance (HDA) capabilities. This limitation necessitated a reliance on mechanical hazard tolerance [1] and the selection of landing sites in terrains deemed exceptionally safe, where 95-99% of the area within the landing ellipse was required to be hazard-free [2]. However, the reliance on statistical estimations of landing safety—predominantly based on limited reconnaissance data from remote sensing—has proven to be inadequate. A notable instance involves NASA’s Viking 1 Mars lander in 1976. Despite thorough planning and examination of the landing site [3], the lander was unexpectedly close to a significant obstacle, a large boulder referred to as Big Joe, posing a potential risk to the mission’s success (Figure 2(a)). The necessity for autonomous HDA was underscored during the Apollo program, which was the first demonstration of real-time HDA via pilot-in-the-loop guidance and navigation. The Apollo 15 mission in 1971, in particular, highlighted the challenges associated with manual HDA [4]; even under near-ideal lighting conditions, navigating and ensuring a safe landing was demanding, resulting in the dramatic tilt of the spacecraft (Fig. 2(b)).
Figure 2: Two representative examples supporting the need for autonomous hazard detection and avoidance (HDA) capability. (a) A boulder, “Big Joe,” about 2 meters long and about 8 meters apart from the Viking 1 Mars lander, posed a potential risk to the mission. (Image ID: PIA00572, JPL Photojournal [5]) (b) The Apollo 15 lunar lander straddling next to a small hazardous crater, resulting in a tilt of $11^{\circ}$; only $1^{\circ}$ away from the tolerance. (Image ID: AS15-87-11818, Apollo 15 Map and Image Library [6])
@@
As the feasibility of autonomous spacecraft landing is increasingly demonstrated, the demand for even more advanced capabilities grows, with the aim of achieving anytime and anywhere global and safe landing capabilities across the solar system—a milestone articulated by NASA’s PL&HA program [24]. A notable challenge in this endeavor is the need for real-time mapping to support active TRN and HDA under any lighting condition during descent [25]. Addressing this requires the development of advanced sensor suites that balance performance with manageable SWaP demands, alongside efficient processing algorithms. This paper introduces a novel approach to stochastic terrain mapping and real-time algorithms that enable efficient terrain analysis and landing safety evaluation with limited data. This method not only potentially reduces the requirements for HD sensors, e.g., range and resolution, but also relaxes their effective operational conditions. Moreover, it offersmat innovative tools for uncertainty quantification in landing safety assessment, paving the way for new stochastic guidance algorithms that can accommodate landing site uncertainty during the final stages of AHDA.
@@ 1.2 Related Works and Preliminaries
The pioneering work of the LiDAR-based HD algorithm by Johnson et al. [15, 16, 41] proposed estimating the base terrain by applying the least median square (LMedsq) fitting. Later, Xiao et al. [42] identified that the LMedsq approach is vulnerable to a large number of outliers and developed a robust estimator for the reference plane by combining the M-estimator sample consensus (MSAC) with a modified iterative Kth ordered scale estimator (IKOSE). In 2019, Xiao et al. applied thin plane spline (TPS) interpolation for computing the base terrain more efficiently, avoiding repeated local plane fittings over different locations [43].
Those HD algorithms with the base terrain approximation do not provide accurate slope and roughness evaluation upon touchdown, as illustrated in Fig. 4. To address this, in the Autonomous Landing and Hazard Avoidance Technology (ALHAT) program, NASA’s Jet Propulsion Laboratory (JPL) proposed a more exact HD algorithm that evaluates deterministic slope and probabilistic roughness safety over the terrain using the landing plane as a reference plane [44]. The improved performance of the proposed algorithm was demonstrated in field tests [45], but the algorithm has proved computationally too expensive, especially for high-resolution digital elevation maps (DEMs) [46].
@@ 2.3 Local GRF Regression via Delaunay Triangulation
Given the triangulation, we infer the terrain topography by the GRF regression to each triangular segment independently. Specifically, we apply Eqs. (5) and (6) to estimate the elevation inside the triangle based on the three vertices. Now, the matrix $\bm{K}^{\epsilon}$ is reduced to a 3x3 size, allowing for the analytical computation of its inverse and accelerating the process. Figure 6 shows the illustration of the local GRF regression via Delaunay triangulation.
Figure 6: Illustration of the local GRF regression via Delaunay triangulation. Once the PCD is triangulated, the mean and variance of the elevations of pixels inside each triangle are estimated with the GRF regression model only based on the three vertices.
@@ 3.2 Baseline HD Algorithm
Figure 7: An example showing the highest local elevation within the lander footprint does not necessarily result in the worst roughness; rock A has a higher elevation but results in smaller roughness than rock B.
Therefore, in our baseline HD algorithm, the probability of roughness safety is evaluated for all pixels within the lander footprint for each orientation across the DEM. Given the number of pixels for a unit length, denoted by $p$, the algorithm’s complexity is $O(p^{5})$, which is still the same as the ALHAT HD algorithm [46].
@@ 3.3 Provably Conservative Real-Time Deterministic HD Algorithm
For an arbitrary lander orientation $\theta$, let $r_{z}(\theta,\bm{\gamma})$ be the signed distance from a point on the terrain, $(\bm{\gamma}=(x,y),z)$, to its projection in z-direction onto the landing plane:
@@ 3.4 Extension to Gaussian DEM Input
$\operatorname{\mathbbm{P}}(\tilde{s}<\bar{s})\approx\Phi\left(\frac{h_{0}\tan(\bar{s})-\mu_{LL}}{\sigma_{LL}}\right),\quad\operatorname{\mathbbm{P}}(\tilde{r}<\bar{r})\approx\Phi\left(\frac{\bar{r}-\mu_{UL}}{\sigma_{UL}}\right)$ (17)
@@ 4.1 Conservativeness Demonstration
First, we demonstrate the conservativeness of the proposed HD algorithm using perfect observations. While we have theoretically proven the conservativeness of the algorithm, we also present numerical demonstrations. Here, we utilize real terrain information, specifically the Mars digital terrain model (DTM) obtained by the High Resolution Imaging Science Experiment (HiRISE) camera.
@@
$c$ Precision Recall
@@
Figure 12: Visualization of safety prediction results across varying terrain complexities. The top row shows the DEM with perfect observation, the second row shows the ground truth of landing safety, and the third row shows the predicted safety maps. The map size is 100 meters by 100 meters in 10 cm/pixel resolution for all maps. White pixels are the safe pixels in the safety maps.
Figure 12 shows visualized results for varying terrain complexities; from left to right, terrain complexity increases. The top row shows the DEM with perfect observation, and the second row shows the ground truth of landing safety. The third row shows the predicted safety maps by the proposed algorithm. We observe that the number of pixels predicted to be safe decreases as terrain complexity increases.
@@ 4.2 Performance Analysis with Simulated Testbed
In practice, obtaining perfect DEMs is challenging, if not impossible. When a planetary lander performs terrain sensing upon arrival at a planetary body, accurately measuring every detail of the surface is difficult. Specifically, we focus on topographic uncertainty due to sparse terrain measurements. In this subsection, we evaluate the performance of the proposed algorithm for such sparse terrain measurements using simulated testbed data.
@@ 4.2.1 Data Generation
Scan Configuration Coverage (m) GSD (m) Noise ($3\sigma_{\text{LiDAR}}$)
Range (m) Angle (°) x y x y (cm)
@@
Table 4: Nominal coverages, ground sample distances (GSDs), and Gaussian noise standard deviations for each LiDAR scan configuration. The presented values are provided for reference purposes only; the simulated scans exhibit variations in coverage and GSDs due to terrain interactions and projection distortions.
@@ 4.2.2 DEM Construction Quality
LiDAR Scan RMSE (m) NLPD
Range (m) Angle (°) Baseline Proposed Baseline Proposed
@@
Table 5: Comparison of baseline and proposed algorithms for DEM reconstruction quality across different LiDAR scan configurations. Results are based on a simulated testbed with 500 rocks, each 1 meter in diameter, placed on a flat 200 x 200 meter surface.
Table 5 presents the numerical results for DEM reconstruction performance in the simulated testbed. Except for the optimal observation condition at the closest range (200 m) with the LiDAR directly pointing downward, the proposed algorithm outperforms the baseline algorithm. Both algorithms show a trend of decreased performance for more challenging LiDAR scan configurations, with less accuracy at longer ranges and larger observational angles.
Figure 14: DEMs constructed by the baseline algorithm over a 10 x 10 meter segment. From top to bottom row, the observational range increases from 200 to 1000 m. From left to right column, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
Figure 15: Mean DEMs constructed by the proposed algorithm over a 10 x 10 meter segment. From top to bottom row, the observational range increases from 200 to 1000 m. From left to right column, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
Figure 16: Variance DEMs constructed by the proposed algorithm over a 10 x 10 meter segment. From top to bottom row, the observational range increases from 200 to 1000 m. From left to right column, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
@@
Figure 17: Comparisons of the true DEM, baseline DEM, and DEM distribution by the proposed algorithm. The left and right columns show results for rocks A and B, as indicated in Fig. 13. Each row shows results for different LiDAR scan configurations as specified in their titles.
@@ 4.2.3 Safety Prediction Performance
Scan Configuration Precision (Slope) Precision (Roughness) Recall (Slope) Recall (Roughness)
Range (m) Angle (°) Baseline Proposed Baseline Proposed Baseline Proposed Baseline Proposed
@@
Table 6 presents the precision and recall for slope and roughness safety across all LiDAR scan configurations. In all cases, the proposed algorithm outperforms the baseline algorithm in terms of precision, indicating that the proposed algorithm provides more reliable safe site predictions than the baseline. This improved reliability comes at the cost of detecting fewer safe sites out of all true safe sites, as indicated by the reduced recall for the proposed algorithm. The benefit of higher precision with the proposed algorithm is more significant for roughness safety than for slope safety and under more challenging observational conditions with larger observational ranges or angles.
@@
Figure 19: Roughness hazard missing rates over a 10 x 10 meter segment for the baseline algorithm. A hazard missing rate of 0 is an accurate prediction, closer to 1 (red) indicates more missed hazards, and closer to -1 (blue) indicates overly conservative hazard assessments. From top to bottom, the observational range increases from 200 to 1000 m. From left to right, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
Figure 20: Roughness hazard missing rates over a 10 x 10 meter segment for the proposed algorithm. A hazard missing rate of 0 is an accurate prediction, closer to 1 (red) indicates more missed hazards, and closer to -1 (blue) indicates overly conservative hazard assessments. From top to bottom, the observational range increases from 200 to 1000 m. From left to right, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
@@
Figure 20 demonstrates that the proposed algorithm can detect roughness hazards even under challenging observational conditions. For observations at a range of 1000 m, the algorithm increases safety uncertainty in response to the large GSDs, resulting in a reduced hazard missing rate.
@@
Figure 22: Slope hazard missing rates over a 10 x 10 meter segment for the baseline algorithm. A hazard missing rate of 0 is an accurate prediction, closer to 1 (red) indicates more missed hazards, and closer to -1 (blue) indicates overly conservative hazard assessments. From top to bottom, the observational range increases from 200 to 1000 m. From left to right, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.
Figure 23: Slope hazard missing rates over a 10 x 10 meter segment for the proposed algorithm. A hazard missing rate of 0 is an accurate prediction, closer to 1 (red) indicates more missed hazards, and closer to -1 (blue) indicates overly conservative hazard assessments. From top to bottom, the observational range increases from 200 to 1000 m. From left to right, the observational angle increases from 0° to 60° from the nadir direction. The black dots denote the LiDAR PCD.

### v2

@@ 1 Introduction
This paper presents a novel real-time stochastic hazard detection algorithm capable of handling topographic uncertainty arising from sparse terrain measurement. First, we develop a novel real-time Gaussian DEM construction algorithm that utilizes Delaunay triangulation and a localized Gaussian Random Field (GRF) model for efficient evaluation. Next, we introduce an innovative real-time stochastic hazard detection algorithm that converts Gaussian DEMs into stochastic safety maps. A fast and provably conservative HD algorithm is developed for standard DEM input, which is then extended to accommodate Gaussian DEM input. Finally, the effectiveness of these proposed methodologies is demonstrated through a combination of real and synthetic terrain data. The proposed pipeline is shown in Fig. 1
@@ 1.1 Background
Historically, the deployment of planetary landers has been constrained by the absence of onboard hazard detection and avoidance (HDA) capabilities. This limitation necessitated a reliance on mechanical hazard tolerance [1] and the selection of landing sites in terrains deemed exceptionally safe, where 95-99% of the area within the landing ellipse was required to be hazard-free [2]. However, the reliance on statistical estimations of landing safety—predominantly based on limited reconnaissance data from remote sensing—has proven to be inadequate. A notable instance involves NASA’s Viking 1 Mars lander in 1976. Despite thorough planning and examination of the landing site [3], the lander was unexpectedly close to a significant obstacle, a large boulder referred to as Big Joe, posing a potential risk to the mission’s success (Fig. 2(a) [4]). The necessity for autonomous HDA was underscored during the Apollo program, which was the first demonstration of real-time HDA via pilot-in-the-loop guidance and navigation. The Apollo 15 mission in 1971, in particular, highlighted the challenges associated with manual HDA [5]; even under near-ideal lighting conditions, navigating and ensuring a safe landing was demanding, resulting in the dramatic tilt of the spacecraft that was only $1^{\circ}$ away from the tolerance (Fig. 2(b) [6]).
Figure 2: (a) A boulder, “Big Joe,” next to the Viking 1 Mars lander [4]. (b) The Apollo 15 lunar lander straddling next to a small crater [6].
@@
As the feasibility of autonomous spacecraft landing is increasingly demonstrated, the demand for even more advanced capabilities grows, with the aim of achieving anytime and anywhere global and safe landing capabilities across the solar system—a milestone articulated by NASA’s PL&HA program [24]. A notable challenge in this endeavor is the need for real-time mapping to support active TRN and HDA under any lighting condition during descent [25]. Addressing this requires the development of advanced sensor suites that balance performance with manageable SWaP demands, alongside efficient processing algorithms. This paper introduces a novel approach to stochastic terrain mapping and real-time algorithms that enable efficient terrain analysis and landing safety evaluation with limited data. This method not only potentially reduces the requirements for HD sensors, e.g., range and resolution, but also relaxes their effective operational conditions. Moreover, it offers innovative tools for uncertainty quantification in landing safety assessment, paving the way for new stochastic guidance algorithms that can accommodate landing site uncertainty during the final stages of AHDA.
@@ 1.2 Related Works and Preliminaries
The pioneering work of the LiDAR-based HD algorithm by Johnson et al. [15, 16, 41] proposed estimating the base terrain by applying the least median square (LMedsq) fitting, instead of the regular least square fitting [42]. Later, Xiao et al. [43] identified that the LMedsq approach is vulnerable to a large number of outliers and developed a robust estimator for the reference plane by combining the M-estimator sample consensus (MSAC) with a modified iterative Kth ordered scale estimator (IKOSE). In 2019, Xiao et al. applied thin plane spline (TPS) interpolation for computing the base terrain more efficiently, avoiding repeated local plane fittings over different locations [44]. Note that for a coarse digital elevation map whose resolution is as large as the lander diameter, we may skip the terrain fitting and take finite difference-based derivatives to estimate the local slope [45].
Those HD algorithms with the base terrain approximation do not provide accurate slope and roughness evaluation upon touchdown, as illustrated in Fig. 4. To address this, in the Autonomous Landing and Hazard Avoidance Technology (ALHAT) program, NASA’s Jet Propulsion Laboratory (JPL) proposed a more exact HD algorithm that evaluates deterministic slope and probabilistic roughness safety over the terrain using the landing plane as a reference plane [46]. The improved performance of the proposed algorithm was demonstrated in field tests [47], but the algorithm has proved computationally too expensive, especially for high-resolution inputs [48].
@@ 2.3 Local GRF Regression via Delaunay Triangulation
Given the triangulation, we infer the terrain topography by the GRF regression to each triangular segment independently. Specifically, we apply Eqs. (5) and (6) to estimate the elevation inside the triangle based on the three vertices. Now, the matrix $\bm{K}^{\epsilon}$ is reduced to a 3x3 size, allowing for the analytical computation of its inverse and accelerating the process. Figure 6 shows the illustration of the local GRF regression via Delaunay triangulation. For example, elevations within the orange triangle in Fig. 6 are estimated using GRF regressions with the elevations of the triangle’s vertices, shown by the red dots.
Figure 6: The mean and variance of elevations are locally interpolated using GRF regressions with the three nearby samples identified by Delaunay triangulation.
@@ 3.2 Baseline HD Algorithm
Figure 7: The highest local elevation does not necessarily correspond to the worst roughness; rock A has a higher elevation but results in smaller roughness than rock B.
Therefore, in our baseline HD algorithm, the probability of roughness safety is evaluated for all pixels within the lander footprint for each orientation across the DEM. Given the number of pixels for a unit length, denoted by $p$, the algorithm’s complexity is $O(p^{5})$, which is still the same as the ALHAT HD algorithm [48]. For the scalability of HD algorithms, see Appendix 5.1 as well.
@@ 3.3 Provably Conservative Real-Time Deterministic HD Algorithm
For an arbitrary lander orientation $\theta$, let $r_{z}(\theta,\bm{\gamma})$ be the signed distance from a point on the terrain, $(\bm{\gamma}=(x,y),z))$, to its projection in z-direction onto the landing plane:
@@ 3.4 Extension to Gaussian DEM Input
$\mathbbm{P}(\tilde{s}<\bar{s})\approx\Phi\left(\frac{h_{0}\tan(\bar{s})-\mu_{LL}}{\sigma_{LL}}\right),\quad\mathbbm{P}(\tilde{r}<\bar{r})\approx\Phi\left(\frac{\bar{r}-\mu_{UL}}{\sigma_{UL}}\right)$ (17)
@@ 4.1 Conservativeness Demonstration
First, we demonstrate the conservativeness of the proposed HD algorithm using perfect observations. While we have theoretically proven the conservativeness of the algorithm, we also present numerical demonstrations. Here, we utilize real terrain information, specifically the Mars digital terrain model (DTM) obtained by the High Resolution Imaging Science Experiment (HiRISE) camera [62].
@@
$c$ Precision $\uparrow$ Recall $\uparrow$
@@
Figure 12: Visualization of safety prediction results across varying terrain complexities. White pixels are the safe pixels in the safety maps.
Figure 12 shows visualized results for varying terrain complexities; from left to right, terrain complexity increases and the map size is 100 meters by 100 meters in 10 cm/pixel resolution for all maps. The top row shows the DEM with perfect observation, and the second row shows the ground truth of landing safety. The third row shows the predicted safety maps by the proposed algorithm. We observe that the number of pixels predicted to be safe decreases as terrain complexity increases.
@@ 4.2 Performance Analysis with Simulated Testbed
In practice, obtaining perfect DEMs is challenging, if not impossible. When a planetary lander performs terrain sensing upon arrival at a planetary body, accurately measuring every detail of the surface is difficult. Specifically, we focus on topographic uncertainty due to sparse terrain measurements. In this subsection, we evaluate the performance of the proposed algorithm for such sparse terrain measurements using simulated testbed data. Here, as for the hyperparameters of Eq. (4), we set $\ell=1.0$ m and compute $\sigma_{f}$ as the standard deviation of the elevations in the point cloud data.
@@ 4.2.1 Data Generation
Table 4: Nominal coverages, ground sample distances (GSDs), and Gaussian noise standard deviations for each LiDAR scan configuration.
Scan Configuration Coverage, m GSD, m Noise, cm
Range, m Angle, deg x y x y $3\sigma_{\text{LiDAR}}$
@@
  • The presented values are provided for reference purposes only; the simulated scans exhibit variations in coverage and GSDs due to terrain interactions and projection distortions.
@@ 4.2.2 DEM Construction Quality
Table 5: Comparison of baseline and proposed algorithms for DEM reconstruction quality across different LiDAR scan configurations.
LiDAR Scan RMSE, m $\downarrow$ NLPD $\downarrow$
Range, m Angle, deg Baseline Proposed Baseline Proposed
@@
Table 5 presents the numerical results for DEM reconstruction performance in the simulated testbed. The results are based on a simulated testbed with 500 rocks, each 1 meter in diameter, placed on a flat 200 x 200 meter surface. Except for the optimal observation condition at the closest range (200 m) with the LiDAR directly pointing downward, the proposed algorithm outperforms the baseline algorithm. Both algorithms show a trend of decreased performance for more challenging LiDAR scan configurations, with less accuracy at longer ranges and larger observational angles.
Figure 14: DEMs constructed by the baseline algorithm over a 10 x 10 meter segment for different ranges and angles of the observation. The black dots denote the LiDAR PCD.
Figure 15: Mean DEMs constructed by the proposed algorithm over a 10 x 10 meter segment for different ranges and angles of the observation. The black dots denote the LiDAR PCD.
Figure 16: Variance DEMs constructed by the proposed algorithm over a 10 x 10 meter segment for different ranges and angles of the observation. The black dots denote the LiDAR PCD.
@@
Figure 17: Comparisons of the true DEM, baseline DEM, and DEM distribution by the proposed algorithm.
@@ 4.2.3 Safety Prediction Performance
Precision $\uparrow$ Recall $\uparrow$
Scan Configuration Slope Roughness Slope Roughness
Range, m Angle, deg Baseline Proposed Baseline Proposed Baseline Proposed Baseline Proposed
@@
Figure 19: Roughness hazard missing rates over a 10 x 10 meter segment for the baseline algorithm. The black dots denote the LiDAR PCD.
Figure 20: Roughness hazard missing rates over a 10 x 10 meter segment for the proposed algorithm. The black dots denote the LiDAR PCD.
@@
Figure 20 demonstrates that the proposed algorithm can detect roughness hazards even under challenging observational conditions. For observations at a range of 1000 m, the algorithm increases safety uncertainty in response to the large GSDs, resulting in a reduced hazard missing rate. Quantitative results for all cases are summarized in Table 6.
@@
Figure 22: Slope hazard missing rates over a 10 x 10 meter segment for the baseline algorithm. The black dots denote the LiDAR PCD.
Figure 23: Slope hazard missing rates over a 10 x 10 meter segment for the proposed algorithm. The black dots denote the LiDAR PCD.
@@
4.3 Performance Analysis with Realistic Data
To evaluate the algorithm’s performance on realistic data, we selected ten distinct locations on the Mars Digital Terrain Model (DTM) for the candidate ExoMars landing site in Hypanis Valles [62]. The visualization of this realistic dataset with region specifications is given in Appendix 5.3.
Here, we also conducted a sensitivity analysis on the hyperparameter $\ell$ of the Gaussian Random Field (GRF) regression described in Eq. (4). The prior standard deviation of elevation $\sigma_{f}$ was computed based on the distribution of the Point Cloud Data (PCD) elevations. Specifically, Table 7 presents the comparison of precision and recall for various $\ell$ values at a range of 500 m with the terrain sensor oriented nadir. The general trend indicates that the proposed algorithm is more conservative than the baseline, which aligns with the results from the simulated testbeds; the proposed algorithm demonstrates higher precision, while the baseline exhibits higher recall. Table 7 further demonstrates that smaller values of $\ell$ result in more conservative estimates, yielding higher precision and lower recall. This behavior is consistent with the physical interpretation of $\ell$; the hyperparameter $\ell$ defines the degree of correlation between nearby elevations, as reflected in the kernel’s formulation in Eq. (4).
This trend is visually demonstrated by Figs. 24 and 25, which illustrate the results for region R4 with $\ell=1.0$ m. In particular, Fig. 25 highlights that the baseline algorithm fails to detect most roughness hazards, whereas the proposed algorithm successfully identifies the majority of them.
Table 7: Comparison of precision and recall for different $\ell$ values at 500m range; $\ell$ in meters.
Slope Roughness
Baseline Proposed Baseline Proposed
$\ell=0.3$ $\ell=1.0$ $\ell=5.0$ $\ell=0.3$ $\ell=1.0$ $\ell=5.0$
Precision $\uparrow$ 0.9515 0.9999 0.9999 0.9997 0.8741 0.9995 0.9906 0.9795
Recall $\uparrow$ 0.9960 0.3417 0.5113 0.6098 0.9999 0.0030 0.0562 0.1389
Figure 24: Performance comparison of baseline and proposed algorithms for realistic terrain: slope case.
Figure 25: Performance comparison of baseline and proposed algorithms for realistic terrain: roughness case.